# NeRF-VO: Real-Time Sparse Visual Odometry with Neural Radiance Fields

### 20 Dec 2023 v1

@@ 5.1 Evaluation Datasets
5.2 Baselines
We compare our method against the SOTA SLAM method NeRF-SLAM [32], which also uses neural implicit representations for mapping. Additionally, where applicable, we report results from the contemporary methods DIM-SLAM [21], NICER-SLAM [58], Orbeez-SLAM [5], GO-SLAM [55] and HI-SLAM [54]. To evaluate the benefits of our back-end, we compare the camera tracking accuracy against DPVO [42] as well as the visual odometry version of DROID-SLAM [41] (DROID-VO), which is used as the tracking module in NeRF-SLAM. These models are evaluated under the same conditions as ours.
5.3 Metrics
We evaluate our method on three tasks: camera tracking, novel view synthesis, and 3D dense reconstruction. For camera tracking, we align the estimated trajectory to the ground truth using the Kabsch-Umeyama [17, 43] algorithm and evaluate the accuracy using the absolute trajectory error (ATE RMSE) [37]. For novel view synthesis, we assess the quality of the rendered RGB images using peak signal-to-noise ratio (PSNR), structural similarity (SSIM) [47], and learned perceptual image patch similarity (LPIPS) [53]. To compute these metrics, we select 125 equally spaced RGB frames along the trajectory. To render the exact view of these evaluation frames, we transform the ground truth camera poses to the coordinate system of the implicit function. For 3D reconstruction, we consider accuracy, completion, and recall. We render the predicted mesh from our neural implicit representation. Analogous to [32, 21, 58, 55, 54] we exclude regions not observed by any camera from the evaluation. We align both meshes using the Iterative Closest Point (ICP) [33] algorithm. To ensure statistical validity, we report the average of 5 independent runs.
5.4 Camera Tracking
Table 1 reports the camera tracking results on the Replica dataset [36, 38], showing that our model achieves the lowest average ATE RMSE over all scenes. The results of DIM-SLAM [21] and NICER-SLAM [58] are taken from their original papers. Other methods have been evaluated by us under exactly the same conditions as ours. While DIM-SLAM [21] and NeRF-SLAM [32] perform better on some scenes, their performance deteriorates on others, whereas our method remains relatively stable across all scenes.
| Model | rm-0 | rm-1 | rm-2 | of-0 | of-1 | o-2 | of-3 | of-4 | Avg. |
@@ 5.4 Camera Tracking
Table 2 reports the tracking accuracy on six commonly tested scenes from ScanNet [8]. Our method outperforms concurrent visual odometry methods whose tracking module only applies implicit “local loop closure” (L-LC) imposed by the scene representations, such as HI-SLAM (VO), DROID-VO, NeRF-SLAM, and DPVO. In general, we can conclude that the full SLAM methods incorporating global loop closure (G-LC) and scene-wide global bundle adjustment perform better on these large scenes.
@@
5.5 3D Reconstruction
Table 4 demonstrates the superior 3D reconstruction performance of our model over all compared concurrent works on the Replica [36] dataset. Our method outperforms SOTA in accuracy, completion, and recall on four scenes and on average. The inferior performance on the office0 and office1 scenes is the result of scene-specific characteristics that complicate monocular depth predictions: these scenes have detail-rich wall paintings that lead to artifacts in monocular depth estimation. We notice that this is also the case for NICER-SLAM [58], which similarly relies on a monocular depth prediction network. A qualitative comparison of the 3D reconstruction is provided in Fig. 3.
@@
Table 5 reports the 3D reconstruction performance of our method and NeRF-SLAM [32] on three real-world datasets: 7-Scenes [35], ScanNet [8], and our custom dataset. The results show that our model outperforms NeRF-SLAM on almost all scenes and on average on all datasets. The scene-level results are available in the supplementary material. Unfortunately, none of the concurrent papers report 3D reconstruction performance on any dataset other than the synthetic Replica [36], thus we are unable to benchmark our model against other methods besides NeRF-SLAM.
@@
5.6 Novel View Synthesis
We evaluated our model for novel view synthesis on Replica [36]. The results can be found in table 6. They indicate mixed performance on the different metrics. While NICER-SLAM performs best on SSIM, NeRF-SLAM and our method are almost on par. NeRF-SLAM performs best in PSNR, yet our approach is almost equal in most scenes. Considering the LPIPS metric, our model shows a significantly superior performance on all scenes, indicating a high subjective similarity for a human observer, underlining the photometric strength of our approach. This can also be observed in the qualitative results in Fig. 4, where the rendered image and ground truth are almost indistinguishable.
5.7 Runtime Analysis
We have conducted a runtime and memory utilization analysis of our method. A comparison with concurrent work is presented in Table 7. The results indicate that our model achieves the highest tracking FPS while leaving the smallest GPU memory footprint. This is primarily due to our runtime- and memory-efficient sparse visual tracking module, and dense geometry enhancement paradigm. In terms of mapping runtime, our method is in the midfield of all competing works. However, it must be emphasized that the two faster methods GO-SLAM [55] and NeRF-SLAM [32] rely on the highly optimized NeRF implementation by Müller et al. [26] using custom CUDA kernels, while our work, similar to DIM-SLAM [21] and NICER-SLAM [58], uses an unoptimized Python implementation, yet our model achieves twice the FPS compared to these Python-based alternatives. Finally, we have also conducted an ablation study on the impact of skipping input frames to reduce runtime, which demonstrates stable and promising performance even with a fourfold speedup (see Sec. 5.8).
@@ 5.7 Runtime Analysis
5.8 Ablations
Dense Depths. We performed an ablation to validate that using aligned dense depths from a monocular estimator does indeed improve performance over directly using the sparse depth patches from our tracking module. The results can be found in Table 8. The substantial performance improvement from using dense depths indicates that optimizing a NeRF with only a few pixel depths per frame is not sufficient to achieve SOTA results.
@@ 5.8 Ablations
| sparse depth | 0.48 | 8.16 | 6.49 | 72.5 |
| dense depth | 0.43 | 2.81 | 3.59 | 86.3 |
Scale Alignment. As discussed in Sec. 4, the sparse depths from our tracking module and the dense depths from the monocular estimator are skewed relative to each other, which requires an alignment strategy. In Table 9 we compare five linear alignment schemes: No alignment of sparse and dense depths (none), alignment based on the minimum and maximum of each depth map (min-max), the least squares alignment proposed by [48] (least-squares), our probability-based alignment method (ours) as well as its relaxed alternative (see Sec. 4). Our choice was confirmed as the probability-based alignment performs best.
| ATE ↓ | Acc. ↓ | Com. ↓ | Rec. ↑ | |
@@
| Total Frames | Speedup | ATE ↓ | Acc. ↓ | Com. ↓ | Rec. ↑ |
| 100% | 1 | 0.43 | 2.81 | 3.59 | 86.3 |
| 50% | 2 | 0.46 | 2.85 | 3.61 | 86.6 |
| 25% | 4 | 0.50 | 2.83 | 3.70 | 85.9 |
| 12.5% | 8 | 3.88 | 6.34 | 6.13 | 77.1 |
Frame Skipping. Analogous to the authors of GO-SLAM [55], we performed a frame skipping ablation to assess what percentage of the total frames is necessary to still achieve SOTA performance. This has a direct impact on the runtime and performance of our model. The results are shown in Table 10. We discovered that, even with 25% of all frames, our approach still achieves nearly equivalent results, with performance degrading only after a reduction to 12.5% (every 8th frame). This means that skipping 75% of the frames could lead to a fourfold speedup of the model, allowing its use in real-time applications.
6 Conclusion
We introduce NeRF-VO, a novel neural visual odometry system that combines a learning-based sparse visual odometry for pose tracking, a monocular-depth prediction network for inferring dense geometry cues, and a specifically designed neural radiance field optimization for pose and dense geometry refinement. NeRF-VO surpasses SOTA methods, demonstrating superior pose estimation accuracy and delivering high-quality dense mapping, all while maintaining low pose tracking latency and GPU memory consumption. As part of future research, it would be compelling to integrate visual or geometric constraints obtained from the neural scene representation into the pose tracking front-end. This integration has the potential to further reduce drift and improve the accuracy of initial pose tracking.

### 16 Jul 2024 v2

@@ V-A Evaluation Datasets
V-B Baselines
We compare our method against the SOTA SLAM method NeRF-SLAM [18], which also uses neural implicit representations for mapping. Additionally, where applicable, we report results from the contemporary methods DIM-SLAM [20], NICER-SLAM [19], Orbeez-SLAM [21], GO-SLAM [22] and HI-SLAM [23]. To evaluate the benefits of our back-end, we compare the camera tracking accuracy against DPVO [29] as well as the visual odometry version of DROID-SLAM [28] (DROID-VO), which is used as the tracking module in NeRF-SLAM.
V-C Metrics
We evaluate our method on three tasks: camera tracking, novel view synthesis, and 3D dense reconstruction. For camera tracking, we align the estimated trajectory to the ground truth using the Kabsch-Umeyama [57, 58] algorithm and evaluate the accuracy using the absolute trajectory error (ATE RMSE) [59]. For novel view synthesis, we assess the quality of the rendered RGB images using peak signal-to-noise ratio (PSNR), structural similarity (SSIM) [60], and learned perceptual image patch similarity (LPIPS) [61]. To compute these metrics, we select 125 equally spaced RGB frames along the trajectory. To render the exact view of these evaluation frames, we transform the ground truth camera poses to the coordinate system of the implicit function. For 3D reconstruction, we consider accuracy, completion, and recall. We render the predicted mesh from our neural implicit representation. Analogous to [18, 20, 19, 22, 23] we exclude regions not observed by any camera from the evaluation. We align both meshes using the Iterative Closest Point (ICP) algorithm. To ensure statistical validity, we report the average of 5 independent runs.
V-D Camera Tracking
Table I reports the camera tracking results on the Replica dataset [1, 2], showing that our model achieves the lowest average ATE RMSE over all scenes. The results of DIM-SLAM [20] and NICER-SLAM [19] are taken from their original papers. While DIM-SLAM [20] and NeRF-SLAM [18] perform better on some scenes, their performance deteriorates on others, whereas our method remains relatively stable across all scenes. Table II reports the tracking accuracy on six commonly tested scenes from ScanNet [52]. Our method outperforms concurrent visual odometry methods whose tracking module only applies implicit “local loop closure” (L-LC) imposed by the scene representations (maps), such as HI-SLAM (VO), NICER-SLAM, DROID-VO, NeRF-SLAM, and DPVO. In general, we can conclude that the full SLAM methods incorporating global loop closure (G-LC) and scene-wide global bundle adjustment perform better on these large scenes. Finally, Table III shows the performance of our method and NeRF-SLAM [18] as well as their respective tracking modules on our custom dataset, demonstrating the superior camera tracking performance of our model.
| Model | rm-0 | rm-1 | rm-2 | of-0 | of-1 | of-2 | of-3 | of-4 | Avg. |
@@ V-D Camera Tracking
V-E 3D Reconstruction
Table IV demonstrates the superior 3D reconstruction performance of our model over all compared concurrent works on the Replica [1] dataset. Our method outperforms SOTA in accuracy, completion, and recall on four scenes and on average. The inferior performance on the office0 and office1 scenes is the result of scene-specific characteristics that complicate monocular depth predictions: these scenes have detail-rich wall paintings that lead to artifacts in monocular depth estimation. We notice that this is also the case for NICER-SLAM [19], which similarly relies on a monocular depth prediction network. A qualitative comparison of the 3D reconstruction is provided in Fig. 2. Table V reports the averaged 3D reconstruction performance of our method and NeRF-SLAM [18] on three real-world datasets: 7-Scenes [54], ScanNet [52], and our custom dataset. The results show that our model outperforms NeRF-SLAM on average on all datasets. Unfortunately, none of the concurrent papers report 3D reconstruction performance on any dataset other than the synthetic Replica [1], thus we are unable to benchmark our model against other methods besides NeRF-SLAM.
@@
V-F Novel View Synthesis
We evaluated our model for novel view synthesis on Replica [1]. The results can be found in Table VI. They indicate mixed performance on the different metrics. While NICER-SLAM performs best on SSIM, NeRF-SLAM and our method are almost on par. NeRF-SLAM performs best in PSNR, yet our approach is almost equal in most scenes. Considering the LPIPS metric, our model shows a significantly superior performance on all scenes, indicating a high subjective similarity for a human observer, underlining the photometric strength of our approach.
V-G Runtime Analysis
We have conducted a runtime and memory utilization analysis of our method. A comparison with concurrent work is presented in Table VII. The results indicate that our model achieves the highest tracking FPS while leaving the smallest GPU memory footprint. This is primarily due to our runtime- and memory-efficient sparse visual tracking module, and dense geometry enhancement paradigm. In terms of mapping runtime, our method is in the midfield of all competing works. However, it must be emphasized that the two faster methods GO-SLAM [22] and NeRF-SLAM [18] rely on the highly optimized NeRF implementation by Müller et al. [6] using custom CUDA kernels, while our work, DIM-SLAM [20] and NICER-SLAM [19] use various unoptimized Python implementations, with our model achieving twice the FPS compared to these Python-based alternatives.
@@ V-G Runtime Analysis
V-H Ablations
Dense Depths. We performed an ablation to validate that using aligned dense depths from a monocular estimator does indeed improve performance over directly using the sparse depth patches from our tracking module. The results can be found in Table VIII. The substantial performance improvement from using dense depths indicates that optimizing a NeRF with only a few pixel depths per frame is not sufficient to achieve SOTA results.
@@ V-H Ablations
| Ours | 0.43 | 2.81 | 3.59 | 86.3 |
| W/O dense depth | 0.48 | 8.16 | 6.49 | 72.5 |
| W/O normals | 0.45 | 3.22 | 3.79 | 83.7 |
Surface Normals. To evaluate the benefit of supervising NeRF optimization with surface normal priors, we performed an ablation by comparing our system to a baseline with the normal consistency loss weight set to zero. The results in Table VIII demonstrate an increase in 3D reconstruction performance by using normal supervision. However, enabling normal supervision slightly increases the runtime, hence one might choose to trade the accuracy gain of normal supervision for a faster runtime.
Scale Alignment. As discussed in Sec. IV, the sparse depths from our tracking module and the dense depths from the monocular estimator are skewed relative to each other, which requires an alignment strategy. In Table IX we compare five linear alignment schemes: No alignment of sparse and dense depths (none), alignment based on the minimum and maximum of each depth map (min-max), the least squares alignment proposed by [62] (least-squares), our probability-based alignment method (ours) as well as its relaxed alternative (see Sec. IV). Our choice was confirmed as the probability-based alignment performs best.
| ATE [cm] ↓ | Acc. [cm] ↓ | Com. [cm] ↓ | Rec. [<5cm %] ↑ | |
@@
VI Conclusion
We introduce NeRF-VO, a novel neural visual odometry system that combines a learning-based sparse visual odometry for pose tracking, a monocular prediction network for inferring dense geometry cues, and a neural radiance field optimization for pose and dense geometry refinement. NeRF-VO surpasses SOTA methods, demonstrating superior pose estimation accuracy and delivering high-quality dense mapping, all while maintaining low pose tracking latency and GPU memory consumption. As part of future research, it would be compelling to integrate visual or geometric constraints obtained from the neural scene representation into the pose tracking front-end. This integration has the potential to further reduce drift and improve the accuracy of initial pose tracking.